OpenStudy (anonymous):

given the following function x>0, on which interval is the function decreasing? f(x)= x/ln(11x)

OpenStudy (dumbcow):

function is decreasing when f'(x) < 0 find f'(x) using quotient rule \[f'(x) = \frac{\ln (11x) - 1}{\ln (11x)^{2}} \lt 0\]

OpenStudy (anonymous):

how did you do that? and would i solve it like a regular equation, in order to get the coordinates for the answer?!

OpenStudy (dumbcow):

i took the derivative, have you learned derivatives yet? otherwise you have to graph function to see where its decreasing anyway like i said , f(x) is decreasing when f'(x) < 0 from there you have to solve for x \[\frac{\ln (11x) -1}{\ln (11x)^{2}} < 0\] \[\ln(11x) -1 < 0\] \[\ln(11x) < 1\] \[11x < e\] \[x < \frac{e}{11}\]

OpenStudy (dumbcow):

Note: function not defined when ln(11x) = 0 --> x = 1/11 given x>0, the decreasing interval is (0,1/11) U (1/11, e/11)